Grade 12Mechanics

An alien ship lies in a crafter if depth R/100 on the surface of the planet of radius R. Planet exists very far from our earth. Alien ship have a special Mechanism that can fire it only once with a velocity which is equal to the escape velocity v from surface. Maximum height from the surface upto which this alien ship can reach on firing is

Aashutosh
8 Years agoGrade 12

1 Answer

Sandeep Bhamoo
8 Years ago
mv²/2 -GMm(100)/99R = -GMM/h. Now mv²/2= GMM/R on putting, the maximum height obtained(h)= 99R. Please do approve my answer if correct